OpenStudy (anonymous):

Laplace transform f(t)=2t+1

OpenStudy (anonymous):

\[\alpha{f(x)}=2\alpha{t} +\alpha {1}\] can't put { } around t & 1... can you finish it?

OpenStudy (anonymous):

using table: http://www.intmath.com/laplace-transformation/table-laplace-transforms.php you'll find: \[2/t ^{2} +1/t\]